# Josephson traveling-wave parametric amplifier based on low-intrinsic-loss coplanar lumped-element waveguide

### v2

@@ Appendix D HARMONIC-BALANCE SIMULATIONS OF CP-JTWPAS
Appendix E MONTE CARLO ANALYSIS OF JTWPA GAIN UNDER JUNCTION CRITICAL-CURRENT VARIATIONS
[…]
@@ Appendix E MONTE CARLO ANALYSIS OF JTWPA GAIN UNDER JUNCTION CRITICAL-CURRENT VARIATIONS
To achieve good phase matching with our windowed modulation schemes, we require a significant impedance modulation of 8% along the transmission line. In a simple implementation, this would demand a 16% variation in stub length (modulating $C_{\mathrm{g}}$) alone. Such large geometrical variations could introduce additional challenges in device design. Instead, we can leverage the characteristic-impedance relation $Z=\sqrt{L/C}$ to distribute the modulation between capacitive and inductive components by simultaneously varying both the stub length and junction area. However, substantial variations in the junction area directly affect the critical current, which could potentially impact JTWPA performance and its resilience against fabrication-induced variations. To investigate this trade-off, we performed Monte Carlo simulations comparing two modulation schemes both with 8% impedance modulation base on a similar design to the Hann-window device given in Appendix A. We compared a single-parameter modulation scheme using stub length (varying $C_{\mathrm{G}}$) variation of 16% alone, and a dual-parameter modulation scheme simultaneously varying both stub length and junction area (varying $L_{\mathrm{J}}$, $C_{\mathrm{J}}$) by 8% in a correlated manner. As shown in Fig. 8, we analyzed gain degradation with harmonic balance simulation, by introducing normally distributed variations in the junction critical current. For each variation level (0% to 10% in 0.2% steps), we performed 5–10 Monte Carlo iterations. In each iteration, we randomly varied the junction parameters according to the specified distribution and optimized the pump power ($-78$ to $-69$ dBm) through binary search to achieve the maximum gain. The gain was evaluated by averaging over the 56-GHz signal band with a 7.427-GHz pump frequency. The small gain difference comes from a slight shift in bandgap frequency due to the different modulation schemes when geometric inductance was considered.
@@
Comparing the two impedance-modulation schemes, besides a slight offset on the average gain between the models, we found no significant difference on how JTWPA gain decreased with increasing junction critical-current variation. Both models showed similar robustness, with gain starting to decrease around 5% variation for our designs, and reaching a drop of 3 dB at roughly 8% variation.
Appendix F SYSTEM-TRANSMISSION CALIBRATION AND NOISE CALCULATION
@@
F.1 Microwave network for the Tukey CP-JTWPA noise and backward gain characterization
@@
Figure 9: Measurement setup used in the noise and backward gain characterization of the Tukey CP-JTWPA. The device is cooled to 12 mK in a dilution refrigerator with microwave switches allowing alternating measurements between the CP-JTWPA and the reference qubit-coupled transmission line. The room temperature equipment configuration shown here is arranged for forward transmission and noise measurements. For backward gain measurements, the VNA is reconfigured to connect to the backward input and output ports while keeping the pump turned on.
@@
The Tukey CP-JTWPA is characterized using a bi-directional microwave circuit that enables measurements in both forward- and backward-transmission directions. A vector network analyzer (VNA; Keysight N5232A) serves as the probe signal source for both transmission and noise measurements. The probe signal is combined with a pump signal generated by a microwave source (Rohde & Schwarz SGS100A) at room temperature via a directional coupler before entering the dilution refrigerator. Along the forward path, signals propagate through a series of cryogenic attenuators, a low-pass filter (RLC Electronics F-30-12.4-R), and a cryogenic double-junction circulator (QCY-G0401202AM) before reaching either the Tukey CP-JTWPA or the qubit-coupled transmission line, as determined by the states of two microwave switches (Radiall R572433000). The forward signals then pass through a microwave diplexer (Marki DPX-1114) installed to dissipate any higher-frequency signals generated by the JTWPA. The forward probe signal subsequently travels through an amplification chain comprising circulators, a filter, and a HEMT amplifier (LNF LNC-4-16B) back to room temperature, where it is further amplified and split by a directional coupler to both the VNA for transmission measurement and a spectrum analyzer (Keysight E4405B) for noise measurement. A band-pass filter was installed to remove the strong pump signal that risk overloading the ADC of the spectrum analyzer. In the backward-gain measurement, the VNA is reconfigured to connect to the backward input and output ports. Note that with a broadband noise source such as a JTWPA, image-noise rejection is necessary for an accurate noise-power characterization if superheterodyne detection is employed.
@@
F.2 System-attenuation calibration using qubit-coupled transmission line
@@
$t=1-e^{i\theta_{\mathrm{F}}}\frac{\xi\Gamma_{1}}{2\Gamma_{2}}\frac{1-\frac{i\Delta}{\Gamma_{2}}}{1+\left(\frac{\Delta}{\Gamma_{2}}\right)^{2}+\frac{\Omega^{2}}{\Gamma_{1}\Gamma_{2}}},$ (F.1)
@@
Fig. 10(a) shows transmission data from a 2D scan over probe power and frequency using a VNA, with data normalized to the near-unity transmission obtained by driving the qubit at high probe power past saturation ($\Omega\gg\Gamma_{1},\Gamma_{2}$). The data was fitted using Eq. (F.1), showing good agreement between experiment and theory. To further check the fit accuracy, we performed an additional power sweep with more averaging at $\omega_{\mathrm{p}}/2\pi=5.5633$ GHz where the transmission minimum occurs. Fig. 10(b) shows this data overlaid with a theoretical curve generated using parameters extracted from the initial 2D fit, demonstrating good agreement without further fitting. The obtained parameters are summarized in Table 2.
@@
System attenuation $A_{\mathrm{sys}}^{\mathrm{cal}}$ $-88.20\pm 0.02$ dB
@@
F.3 Translation of system calibration
The noise measurement requires precise knowledge of the system gain from the CP-JTWPA output to the room-temperature digitizer. As the qubit-based calibration was performed using a different sample package, the calibration needs to be translated between the two configurations. The measurement setup uses two microwave switches to alternately connect either the CP-JTWPA or the qubit-coupled line to the measurement path (Fig. 11). Both samples connect to each side of the switches through cables with near-identical transmissions. Package and CP-JTWPA losses were characterized over multiple cooldowns using a 50-Ohm union as reference.
@@ F.3 Translation of system calibration
The system gain $G_{\mathrm{sys}}^{\mathrm{cal}}$ is calculated from the overall system transmission $T_{\mathrm{sys}}$ by accounting for the calibrated system attenuation $A_{\mathrm{sys}}^{\mathrm{cal}}$, the qubit package loss $A_{Q}$, and the CP-JTWPA insertion loss $A_{\mathrm{JTWPA}}$:
$G_{\mathrm{sys}}^{\mathrm{cal}}=T_{\mathrm{sys}}-(A_{\mathrm{sys}}^{\mathrm{cal}}-A_{Q}/2+A_{\mathrm{JTWPA}}),$ (F.2)
where $A_{Q}=-0.57$ dB represents the loss through the qubit package with the qubit coupling point at the middle of the path, and $A_{\mathrm{JTWPA}}=-0.29$ dB is the total insertion loss of the Tukey CP-JTWPA including its package. The system transmission $T_{\mathrm{sys}}$ is measured as the difference in the set power between a probe signal source (VNA in this case) and the received power at the spectrum analyzer which is also used in noise measurement. This calibration procedure yields $G_{\mathrm{sys}}^{\mathrm{cal}}=56.92$ dB.
@@
Figure 11: Measurement setup for system calibration. Two microwave switches enable alternate measurements between the CP-JTWPA path and qubit-coupled-line path, where paths of the same color share near-identical transmissions. System parameters labeled include: calibrated system attenuation $A_{\mathrm{sys}}^{\mathrm{cal}}$, calibrated system gain $G_{\mathrm{sys}}^{\mathrm{cal}}$, qubit package loss $A_{Q}$, and CP-JTWPA insertion loss $A_{\mathrm{JTWPA}}$.
F.4 Noise calculation
@@
Using the calibrated system gain $G_{\mathrm{sys}}^{\mathrm{cal}}$, we measured the noise performance of the CP-JTWPA as illustrated in Fig. 12. We determined the noise power at room temperature in units of photon flux per hertz, $N_{\mathrm{RT}}=P_{\mathrm{SA}}/\mathrm{B}/\hbar\omega_{\mathrm{p}}$, where $P_{\mathrm{SA}}$ is the average noise power measured by a spectrum analyzer, B = 100 Hz is the measurement bandwidth, and $\hbar\omega_{\mathrm{p}}$ is the energy of a photon at $5.5633~{}\mathrm{GHz}$ where the noise measurement was performed. Given the low insertion loss of our CP-JTWPA, we modeled the device as a single input attenuator with attenuation $A_{\mathrm{JTWPA}}$ followed by an ideal parametric amplifier with gain $G_{\mathrm{JTWPA}}$. Through cryogenic attenuation and thermal anchoring of the input line, we assume the input noise remains at the vacuum fluctuation level.
Figure 12: Noise propagation schematic in our measurement setup. The input signal first experiences the CP-JTWPA’s intrinsic loss $A_{\mathrm{JTWPA}}$ before undergoing parametric amplification $G_{\mathrm{JTWPA}}$. The signal then propagates through a calibrated room-temperature amplification chain $G_{\mathrm{sys}}^{\mathrm{cal}}$ before being measured by the spectrum analyzer at room temperature to give $N_{\mathrm{RT}}$. The dashed box indicates the system boundary for calculating total system noise $N_{\mathrm{sys}}$.
@@
To ensure measurement accuracy and account for system noise fluctuations, we performed 100 repeated measurements with interleaved CP-JTWPA pump-on and pump-off states, each lasting approximately one second. This interleaved measurement scheme minimizes the impact of amplifier chain noise drift on the added noise measurement.
When the CP-JTWPA is off, the measured noise at room temperature $N_{\mathrm{RT}}^{\mathrm{off}}$ consists of amplified vacuum noise plus the system added noise:
$N_{\mathrm{RT}}^{\mathrm{off}}=G_{\mathrm{sys}}^{\mathrm{cal}}(N_{\mathrm{vac}}+N_{\mathrm{sys}}^{\mathrm{cal}})$ (F.3)
@@
When the CP-JTWPA is turned on, it contributes both parametric gain $G_{\mathrm{J}}$ and its own added noise $N_{\mathrm{J}}$ after the initial attenuation. The measured noise becomes:
$N_{\mathrm{RT}}^{\mathrm{on}}=G_{\mathrm{sys}}^{\mathrm{cal}}\left(G_{\mathrm{J}}(N_{\mathrm{vac}}+N_{\mathrm{J}})+N_{\mathrm{sys}}^{\mathrm{cal}}\right)$ (F.4)
The difference between these two states directly relates to the CP-JTWPA’s added noise:
$N_{\mathrm{RT}}^{\mathrm{on}}-N_{\mathrm{RT}}^{\mathrm{off}}=G_{\mathrm{sys}}^{\mathrm{cal}}\left(G_{\mathrm{J}}(N_{\mathrm{vac}}+N_{\mathrm{J}})-N_{\mathrm{vac}}\right)$ (F.5)
From this, we extract the added noise of the CP-JTWPA:
$N_{\mathrm{J}}=\frac{(N_{\mathrm{RT}}^{\mathrm{on}}-N_{\mathrm{RT}}^{\mathrm{off}})/G_{\mathrm{sys}}^{\mathrm{cal}}+N_{\mathrm{vac}}}{G_{\mathrm{J}}}-N_{\mathrm{vac}}$ (F.6)
@@
The overall system noise in photo flux per Hertz $N_{\mathrm{sys}}=N_{\mathrm{RT}}/G_{\mathrm{sys}}$ is calculated for both on and off states by treating the CP-JTWPA and subsequent amplification chain (indicated by the dashed box in Fig. 12) as a single amplification system given by $G_{\mathrm{sys}}$. The noise temperature is converted from the system noise as $T_{\mathrm{N}}=N_{\mathrm{sys}}\hbar\omega_{\mathrm{p}}/k_{\mathrm{B}}$.
Appendix G FABRICATION RECIPE

### v3

@@ E.2 Simulated gain profiles of boxcar CP-JTWPA under impedance mismatch
Appendix F MONTE CARLO ANALYSIS OF JTWPA GAIN UNDER JUNCTION CRITICAL-CURRENT VARIATIONS
Figure 9: Performance comparison of capacitive-only ($C_{\mathrm{G}}$) and capacitive–inductive ($C_{\mathrm{g}}$$L_{\mathrm{J}}$) modulated JTWPAs. The frequency-averaged maximum gain is plotted against the junction critical-current variation in standard-deviation percentage. Individual simulation points (dots) and their moving averages over twenty successive points (solid lines) show gain response over 0–10% variation, showing a similar gain degradation in both modulation schemes.
Our devices implement the dual-parameter modulation scheme described in Section II and Appendix A, where both junction area and stub length are varied in a correlated manner to achieve the required 8% impedance modulation. While this approach reduces the required stub-length variation from 16% to 8% compared to capacitive-only modulation, it intentionally introduces variations in junction critical current along the transmission line. Since uniform junction parameters are typically desired in JTWPAs to minimize performance degradation from fabrication variations, we performed Monte Carlo simulations to verify that this intentional non-uniformity does not compromise the device’s robustness against additional fabrication-induced junction variations.
We compared two approaches, both achieving 8% impedance modulation based on a design similar to the Hann-window device given in Appendix A. The first uses single-parameter modulation with 16% stub-length variation (varying $C_{\mathrm{G}}$) alone, while the second implements our dual-parameter modulation scheme simultaneously varying both stub length and junction area (varying $C_{\mathrm{g}}$, $L_{\mathrm{J}}$) by 8% in a correlated manner.
@@ Appendix F MONTE CARLO ANALYSIS OF JTWPA GAIN UNDER JUNCTION CRITICAL-CURRENT VARIATIONS
As shown in Fig. 9, we analyzed gain degradation through harmonic-balance simulations by introducing normally distributed variations in the junction critical current. For each variation level (0% to 10% in 0.2% steps), we performed 5–10 Monte Carlo iterations. In each iteration, we randomly varied the junction parameters according to the specified distribution and optimized the pump power (78-78 to 69-69 dBm) through binary search to achieve the maximum gain. The gain was evaluated by averaging over the 5–6-GHz signal band with a 7.427-GHz pump frequency. The small gain difference between the two schemes arises from a slight shift in band-gap frequency due to the different modulation approaches when geometric inductance is considered.
@@
Comparing the two impedance-modulation schemes, besides a slight offset in the average gain between the models, we found no significant difference in how JTWPA gain decreased with increasing junction critical-current variation. Both models showed similar robustness, with gain starting to decrease around 5% variation for our designs, and reaching a 3-dB drop at roughly 8% variation.
Appendix G SYSTEM-TRANSMISSION CALIBRATION AND NOISE CALCULATION
@@
G.1 Microwave network for the Tukey CP-JTWPA noise and backward gain characterization
@@
Figure 10: Measurement setup used in the noise and backward gain characterization of the CP-JTWPA. The device is cooled to 12 mK in a dilution refrigerator with microwave switches allowing alternating measurements between the JTWPA and the reference qubit-coupled transmission line. The room temperature equipment configuration shown here is arranged for forward transmission and noise measurements. For backward gain measurements, the VNA is reconfigured to connect to the backward input and output ports while keeping the pump turned on.
@@
The Tukey CP-JTWPA is characterized using a bi-directional microwave circuit that enables measurements in both forward- and backward-transmission directions. A vector network analyzer (VNA; Keysight N5232A) serves as the probe signal source for both transmission and noise measurements. The probe signal is combined with a pump signal generated by a microwave source (Rohde & Schwarz SGS100A) at room temperature via a directional coupler before entering the dilution refrigerator. Along the forward path, signals propagate through a series of cryogenic attenuators, a low-pass filter (RLC Electronics F-30-12.4-R), and a cryogenic double-junction circulator (QCY-G0401202AM) before reaching either the CP-JTWPA or the qubit-coupled transmission line, as determined by the states of two microwave switches (Radiall R572433000). The forward signals then pass through a microwave diplexer (Marki DPX-1114) installed to dissipate any higher-frequency signals generated by the JTWPA. The forward probe signal subsequently travels through an amplification chain comprising circulators, a filter, and a HEMT amplifier (LNF LNC-4-16B) back to room temperature, where it is further amplified and split by a directional coupler to both the VNA for transmission measurement and a spectrum analyzer (Keysight E4405B) for noise measurement. A band-pass filter was installed to remove the strong pump signal that risk overloading the ADC of the spectrum analyzer. In the backward-gain measurement, the VNA is reconfigured to connect to the backward input and output ports. Note that with a broadband noise source such as a JTWPA, image-noise rejection is necessary for an accurate noise-power characterization if superheterodyne detection is employed.
@@
G.2 System-attenuation calibration using qubit-coupled transmission line
@@
$t=1-e^{i\theta_{\mathrm{F}}}\frac{\xi\Gamma_{1}}{2\Gamma_{2}}\frac{1-\frac{i\Delta}{\Gamma_{2}}}{1+\left(\frac{\Delta}{\Gamma_{2}}\right)^{2}+\frac{\Omega^{2}}{\Gamma_{1}\Gamma_{2}}},$ (G.1)
@@
Fig. 11(a) shows transmission data from a 2D scan over probe power and frequency using a VNA, with data normalized to the near-unity transmission obtained by driving the qubit at high probe power past saturation ($\Omega\gg\Gamma_{1},\Gamma_{2}$). The data was fitted using Eq. (G.1), showing good agreement between experiment and theory. To further check the fit accuracy, we performed an additional power sweep with more averaging at $\omega_{\mathrm{p}}/2\pi=5.5633$ GHz where the transmission minimum occurs. Fig. 11(b) shows this data overlaid with a theoretical curve generated using parameters extracted from the initial 2D fit, demonstrating good agreement without further fitting. The obtained parameters are summarized in Table 2.
@@
System attenuation $A_{\mathrm{sys}}^{\mathrm{cal}}$ $88.20\pm 0.02$ dB
@@
G.3 Translation of system calibration
The noise measurement requires precise knowledge of the system gain from the Tukey CP-JTWPA output to the room-temperature digitizer. Since the qubit-based calibration used a separate sample package, the calibration must be translated between the two configurations.
The measurement setup uses two microwave switches to alternately connect either the CP-JTWPA or the qubit-coupled line to the measurement path (see Fig. 12). Each measurement path employs matched cable pairs, with near-identical cables connecting each sample to the input and output switches, respectively, to the minimize transmission difference between the two measurement configurations. We characterized the package losses for both devices in separate cool downs using a 50-$\Omega$ union as reference, then incorporated these losses by assuming symmetric loss distribution.
The system gain $G_{\mathrm{sys}}^{\mathrm{cal}}$ from the CP-JTWPA package output port to the room-temperature spectrum analyzer is calculated from the measured system transmission $|S_{\mathrm{21,sys}}|^{2}$ between the VNA source and spectrum analyzer:
$G_{\mathrm{sys}}^{\mathrm{cal}}=\left[|S_{\mathrm{21,sys}}|^{2}+(A_{\mathrm{sys}}^{\mathrm{cal}}-A_{\mathrm{Qubit}}/2+A_{\mathrm{JTWPA}})\right]\,\,\mathrm{dB},$ (G.2)
@@ G.3 Translation of system calibration
where $A_{\mathrm{sys}}^{\mathrm{cal}}=88.20$ dB is the calibrated system attenuation obtained from the qubit-coupled transmission line, $A_{\mathrm{Qubit}}=0.57$ dB is the qubit package loss, and $A_{\mathrm{JTWPA}}=0.29$ dB is the total CP-JTWPA insertion loss including its package. This calculation yields $G_{\mathrm{sys}}^{\mathrm{cal}}=56.9$ dB.
For calculating the JTWPA added noise, we model package losses as lumped attenuators at the input and output of an otherwise lossless JTWPA. While this approach can overestimate added noise compared to distributed noise models used in previous works [9, 11], the difference is negligible given our low JTWPA and package insertion losses. To properly account for the package loss in our noise calculation, we shift the reference plane of system gain from the JTWPA package output port to the center of the JTWPA package. This yields the device-referenced system gain $\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}=G_{\mathrm{sys}}^{\mathrm{cal}}/(A_{\mathrm{JTWPA}}/2)=56.76$ dB.
@@
Figure 12: Measurement setup for system calibration. Two microwave switches enable alternate measurements between the Tukey CP-JTWPA path and qubit-coupled-line path, where paths of the same color share near-identical transmissions. System parameters labeled include: calibrated system attenuation $A_{\mathrm{sys}}^{\mathrm{cal}}$, calibrated system gain $G_{\mathrm{sys}}^{\mathrm{cal}}$, qubit package loss $A_{\mathrm{Qubit}}$, and CP-JTWPA insertion loss $A_{\mathrm{JTWPA}}$. The device-referenced system gain $\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}$ is obtained by adjusting $G_{\mathrm{sys}}^{\mathrm{cal}}$ for half of the CP-JTWPA package loss to reference the gain to the center of the JTWPA device.
G.4 Noise calculation
@@
Using $\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}$, we calculate the added noise of the CP-JTWPA as illustrated in Fig. 13. We determine the noise power at room temperature in units of photon flux per hertz, $N_{\mathrm{RT}}=P_{\mathrm{SA}}/\mathrm{B}/\hbar\omega_{\mathrm{p}}$, where $P_{\mathrm{SA}}$ is the average noise power measured by the spectrum analyzer, $\mathrm{B}=100$ Hz is the measurement bandwidth, and $\hbar\omega_{\mathrm{p}}$ is the photon energy at the measurement frequency of $5.5633$ GHz. We assume the input noise remains at the vacuum fluctuation level due to cryogenic attenuation and thermal anchoring of the input line.
@@
To ensure measurement accuracy and account for system noise fluctuations, we performed 100 repeated measurements with interleaved CP-JTWPA pump-on and pump-off states, each lasting approximately one second. This interleaved measurement scheme minimizes the impact of amplifier chain noise drift on the added noise measurement. When the CP-JTWPA is off, the measured noise at room temperature $N_{\mathrm{RT}}^{\mathrm{off}}$ consists of amplified vacuum noise $N_{\mathrm{vac}}$ and system added noise $\widetilde{N}_{\mathrm{sys}}^{\mathrm{cal}}$ (at JTWPA reference plane):
$N_{\mathrm{RT}}^{\mathrm{off}}=\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}(N_{\mathrm{vac}}+\widetilde{N}_{\mathrm{sys}}^{\mathrm{cal}})$ (G.3)
@@
When the CP-JTWPA is turned on, it contributes both parametric gain $G_{\mathrm{JTWPA}}$ and its own added noise $N_{\mathrm{JTWPA}}$ to the signal. The measured noise becomes:
$N_{\mathrm{RT}}^{\mathrm{on}}=\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}\left(G_{\mathrm{JTWPA}}(N_{\mathrm{vac}}+N_{\mathrm{JTWPA}})+\widetilde{N}_{\mathrm{sys}}^{\mathrm{cal}}\right)$ (G.4)
The difference between these two states isolates the CP-JTWPA’s contribution:
$N_{\mathrm{RT}}^{\mathrm{on}}-N_{\mathrm{RT}}^{\mathrm{off}}=\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}\left(G_{\mathrm{JTWPA}}(N_{\mathrm{vac}}+N_{\mathrm{JTWPA}})-N_{\mathrm{vac}}\right)$ (G.5)
Solving for the CP-JTWPA added noise yields:
$N_{\mathrm{JTWPA}}=\frac{(N_{\mathrm{RT}}^{\mathrm{on}}-N_{\mathrm{RT}}^{\mathrm{off}})/\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}+N_{\mathrm{vac}}}{G_{\mathrm{JTWPA}}}-N_{\mathrm{vac}}$ (G.6)
Figure 13: Noise-propagation schematic in our measurement setup. The input signal experiences half of the CP-JTWPA package loss ($A_{\mathrm{JTWPA}}/2$), then undergoes parametric amplification ($G_{\mathrm{JTWPA}}$) where the JTWPA added noise ($N_{\mathrm{JTWPA}}$) is introduced, followed by the remaining half of the package loss after. The signal then propagates through the amplification chain ($G_{\mathrm{sys}}^{\mathrm{cal}}$) before being measured by the spectrum analyzer to yield $N_{\mathrm{RT}}$. The blue dashed box indicates the system boundary for calculating the input-referred system noise $N_{\mathrm{sys}}$. The black dashed line indicates the reference plane for the CP-JTWPA added noise calculation, located at the center of the CP-JTWPA package.
@@
The input-referred system noise in photon flux per hertz $N_{\mathrm{sys}}=N_{\mathrm{RT}}/G_{\mathrm{sys}}$ is calculated for both on and off states by treating the CP-JTWPA and subsequent amplification chain (indicated by the dashed box in Fig. 13) as a single amplification system, with a system gain $G_{\mathrm{sys}}=G_{\mathrm{sys}}^{\mathrm{cal}}/A_{\mathrm{JTWPA}}$. The noise temperature is converted from the system noise quanta as $T_{\mathrm{N}}=N_{\mathrm{sys}}\hbar\omega_{\mathrm{p}}/k_{\mathrm{B}}$.
Our uncertainty analysis accounts for both statistical and systematic contributions. Statistical uncertainties are calculated as standard deviations from repeated measurements. For systematic uncertainties, we estimate a $\pm$0.5-dB uncertainty in $\widetilde{G}_{\mathrm{sys}}^{\mathrm{cal}}$, based on typical variations observed in our cryogenic measurements, including cable differences and package-loss variations. To incorporate this systematic uncertainty, we calculate the added noise and system noise at both the upper and lower bounds of this $\pm$0.5-dB range, creating an uncertainty envelope. We then report the total uncertainty as half the width of this envelope, representing the maximum deviation from our central estimate.
Appendix H FABRICATION RECIPE